kq/(x+a)^2-kq/(x-a)^2
1
Expert's answer
2011-03-17T08:06:53-0400
kq/(x+a)^2 - kq/(x-a)^2 =
= (kq * (x-a)^2)/((x+a)^2 * (x-a)^2) - (kq * (x+a)^2)/((x-a)^2 * (x+a)^2) =
= (kq * (x^2-2*x*a+a^2))/((x+a)^2 * (x-a)^2) - (kq * (x^2+2*x*a+a^2))/((x-a)^2 * (x+a)^2) =
= (kq * (x^2-2*x*a+a^2))/((x^2-a^2)^2) - (kq * (x^2+2*x*a+a^2))/((x^2-a^2)^2) =
= kq * (- 4*x*a)/((x^2-a^2)^2)